OpenStudy (anonymous):

A recipe requires 3.5 teaspoons of sugar to make a tart. Which equation shows the number of teaspoons of sugar, y, needed to make x tarts?

OpenStudy (anonymous):

x = 3.5y y = 3.5x y = 3.5 + x x = 3.5 + y

Vocaloid (vocaloid):

hint: number of teaspoons = 3.5*number of tarts

OpenStudy (anonymous):

A

Vocaloid (vocaloid):

not quite, check carefully y = number of teaspoons x = number of tarts

OpenStudy (anonymous):

oh B

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

i see it now thank you c: